OpenStudy (anonymous):

how to solve this y + 3 = 2x -x + y + 2 by using slope intercept ...?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

I'm not sure what that means.. Do you have two equations there?

OpenStudy (anonymous):

its system linear equations

OpenStudy (anonymous):

Well that second line doesn't have an equal sign, so I'm not sure that it's an equation.

OpenStudy (anonymous):

i dont know how to solve that by using slope intercept ..pls help me

OpenStudy (anonymous):

pls. explain ,,? why do you said that it is not equation

OpenStudy (anonymous):

-x + y + 2

OpenStudy (anonymous):

anu poh yan.?

OpenStudy (anonymous):

what is that.?

OpenStudy (anonymous):

he actually pressed shift while typing = sign...so the last + is an = sign :)

OpenStudy (anonymous):

i dont understand :( the problem

OpenStudy (anonymous):

is the equation is -x+y=2 ?? if yes, the answer will be y=x+2 in fist equation we get x+2+3=2x ==> x=5 and y=7